OpenStudy (anonymous):

What is the length of XY( the minor arc? Use your calculator and round your final answer to the nearest hundredth

OpenStudy (anonymous):

@jim_thompson5910 @mathstudent55 @Mertsj @UnkleRhaukus @yummydum

OpenStudy (anonymous):

@mathstudent55 @Mertsj

jimthompson5910 (jim_thompson5910):

XY = (40/360)*(Circumference)

OpenStudy (anonymous):

I got 81?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get for the circumference?

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

wait how do u find the circumference?

OpenStudy (anonymous):

@jim_thompson5910 @Mertsj

OpenStudy (mertsj):

Well, I see that the angle is 40 degrees. I'm trying to think what part of the entire circle that is. How many degrees are in the whole circle?

jimthompson5910 (jim_thompson5910):

C = 2*pi*r C = 2*3.14*9 C = ???

OpenStudy (anonymous):

I got 508.98?

OpenStudy (anonymous):

as my answer @jim_thompson5910 @Mertsj

OpenStudy (mertsj):

That seems like a big answer. How did you get it?

OpenStudy (anonymous):

I did: 360/40= 4 then I did: 2 x 3.14 x 9= 56.52 then I multiplied them both (4 x 56.52) and I got 508.98

OpenStudy (anonymous):

@Mertsj @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sry the site isn't working for me, is it working for you?

jimthompson5910 (jim_thompson5910):

well now it is, but its off and on...

OpenStudy (anonymous):

now it finally is! Lol yeah

OpenStudy (anonymous):

it's fine! I'm having the same issue

jimthompson5910 (jim_thompson5910):

circumference is 56.52

jimthompson5910 (jim_thompson5910):

so XY = (40/360)*(Circumference) XY = (40/360)*(56.52) XY = ???

OpenStudy (anonymous):

actually for 360/40 i got 9 :) I got 508.68

jimthompson5910 (jim_thompson5910):

you're dividing the wrong way

jimthompson5910 (jim_thompson5910):

it's not 360/40 it should be 40/360

OpenStudy (anonymous):

OH that made SUCH a big difference..I didn't think that mattered! Haha

OpenStudy (anonymous):

oops xD Lol I got 6.28 as my answer?

jimthompson5910 (jim_thompson5910):

you can only flip the terms if you are either adding or multiplying

jimthompson5910 (jim_thompson5910):

not dividing or subtracting

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (anonymous):

Ohhh okay got it thank you so much!! :)